# McKean-Vlasov equations and nonlinear Fokker-Planck equations with critical singular Lorentz kernels

### v1

@@ 2.3. Sobolev regularity estimate
$\|\nabla u\|_{L^{\mathfrak{q}}_{t}L_{x}^{\mathfrak{l},1}}\leqslant C\left(\|\nabla u\|_{L^{\mathfrak{q}_{1}}_{t}L_{x}^{\mathfrak{l}_{1}}}+\|\nabla u\|_{L^{\mathfrak{q}_{2}}_{t}L_{x}^{\mathfrak{l}_{2}}}\right)\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|},$
@@
$\|b\cdot\nabla u\|_{L^{\mathfrak{q}}_{t}L_{x}^{\mathfrak{p}}}\leqslant C\|b\|_{L^{\infty}_{t}L_{x}^{2,\infty}}\|\nabla u\|_{L^{\mathfrak{q}}_{t}L_{x}^{\mathfrak{l},1}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|},$
@@
$\|u\|_{L^{\mathfrak{q}}_{t}W^{2,\mathfrak{p}}_{x}}+\|\partial_{t}u\|_{L^{\mathfrak{q}}_{t}L_{x}^{\mathfrak{p}}}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|cu+f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leqslant C\|u\|_{L_{t,x}^{\infty}}\|c\|_{L^{\infty}_{t}(L^{1}_{x}\cap L^{\infty}_{x})}+C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leqslant C{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|f\right|\kern-1.07639pt\right|\kern-1.07639pt\right|},$
@@ 3.1. Existence
[…]
@@ 3.2. Uniqueness
[…]
@@
$-\langle\zeta,v(0)\rangle=\int_{0}^{T}\langle\rho(s),f(s)\rangle\text{\rm{d}}s,\quad\forall f\in C_{c}^{\infty}((0,T)\times{\mathbb{R}}^{d}),$
@@ 4.1. SDEs with supercritical drifts
Let $g$ be a non-negative smooth function on $[0,\infty)$ such that $g^{\prime}(r)\geqslant 0$ for all $r\geqslant 0$, $g(r)=0$ if $20\leqslant r\leqslant 1/$ 2 and $g(r)=1$ if $r>1$. Let
@@
and for $11\leqslant i\leqslant d-$ 1,
$b_{i}(x):=$ $-Nx_{i}r^{1-d}\partial_{x_{d}}H(x)$ (4.3)
@@
for all $x_{d}>0$ and $11\leqslant i\leqslant d-$ 1. Taking into account (4.4) and the anti-symmetry of $b$ about the $x_{d}$-axis, we obtain that $\mathord{{\rm div}}b(x)=0$ if $x\neq 0$.
@@
Since $20<1/(1+\alpha)<1/$ 2, we can choose $N\gg 1$ such that
@@
The above inclusion indicates that the solution paths starting from $x\in{\mathcal{C}}_{\kappa,2}$ stay in the larger cone ${\mathcal{C}}_{1}$ over a certain time interval $[0,t_{x}]$, and return back to the small cone ${\mathcal{C}}_{\kappa}$ at time $t_{x}$.
@@
$V_{t}^{x}(\omega)-x_{d}=&\int_{0}^{t}b_{d}(\omega_{s})\text{\rm{d}}s=N(d-1)\int_{0}^{t}[V^{x}_{s}(\omega)+(B_{s}(\omega))_{d}]^{-\alpha}\text{\rm{d}}s\geqslant 0,\\ \widehat{\omega}_{t}-\hat{x}=&N\alpha\int_{0}^{t}[V^{x}_{s}(\omega)+(B_{s}(\omega))_{d}]^{-\alpha-1}\widehat{\omega}_{s}\text{\rm{d}}s+\widehat{B}_{t}(\omega).\end{aligned}$ (4.13)
@@
for all $x\in\overline{{\mathcal{C}}}_{\kappa,2}\backslash\{0\}$, $\omega\in\Omega_{N}^{x}$ and $t\in[0,t_{x}\wedge\tau(\omega)]$. By virtue of Chaplygin’s Lemma, we derive
@@
for all $x\in\overline{{\mathcal{C}}}_{\kappa,2}\backslash\{0\},\omega\in\Omega_{N}^{x}$ and $t\in[0,t_{x}\wedge\tau(\omega)]$. This along with (4.14) yields that
@@
for all $x\in\overline{{\mathcal{C}}}_{\kappa,2}\backslash\{0\},\omega\in\Omega_{N}^{x}$ and $t\in[0,t_{x}\wedge\tau(\omega)]$.
@@
Moreover, via (4.15), for sufficiently large $N$, $x\in\overline{{\mathcal{C}}}_{\kappa,2}\backslash\{0\}$ and $\omega\in\Omega_{N}^{x}$, we also have
@@
Then $S(\omega)<\overline{\tau}(\omega)<1$, $\omega_{S}\in\partial{\mathcal{C}}_{\kappa,2}\backslash\{0\}$ and $t_{\omega_{S}}<N^{-1}2^{3}<9$. Therefore, the inclusion in (4.11) for $x\in\overline{{\mathcal{C}}}_{\kappa,2}\backslash\{0\}$ and $\omega\in\Omega_{N}^{x}$ also holds, respectively, for the new starting point $\omega_{S}$ and the new path $\theta_{S}\omega$. Thus, we have
@@ 4.2. FPE with supercritical drifts
For this purpose, let $\sigma=\inf\{t>0:\omega_{t}=0\}$. Because the drift $b$ is smooth in ${\mathbb{R}}^{d}\backslash\{0\}$ and bounded on $B_{\varepsilon}^{c}(0)$ for any $\varepsilon>0$, by the standard SDE well-posedness theory, for any $x\neq 0$ away from the origin, the restriction ${\mathbb{P}}_{x}\restriction_{{\mathcal{F}}_{\sigma}}$ of the martingale solution ${\mathbb{P}}_{x}$ to the filtration ${\mathcal{F}}_{\sigma}$ is uniquely determined. Then, for any $x\neq 0$ and any bounded measurable function $f$,
@@
where the last identity was due to the uniqueness assumption and the fact that the regular conditional probability ${\mathbb{P}}_{x}(\cdot|{\mathcal{F}}_{\sigma})(\omega^{\prime})$ is also a martingale solution to (SDE) starting from 00, for ${\mathbb{P}}_{x}$-a.s. $\omega^{\prime}$. Taking into account that the distribution of $(\sigma,\omega_{\cdot\wedge\sigma})$ is uniquely determined under ${\mathbb{P}}_{x}$, we thus infer that any one-dimensional marginal distribution of weak solutions to (SDE) satisfies the same identity (4.2), and thus, is unique for all $x\in{\mathbb{R}}^{d}$.
@@ 5.1. Existence
$\text{\rm{d}}X_{t}^{n}&=K^{n}(t,X_{t}^{n}-y)\rho^{n}(t)(\text{\rm{d}}y)+\sqrt{2}\text{\rm{d}}W_{t}\\ \rho^{n}(t)&=\mathrm{law}(X_{t}^{n}),~{}\rho^{n}(0)=\zeta\in{\mathcal{P}}({\mathbb{R}}^{d}).\end{aligned}\right.$
@@
[…]
@@ 5.2. Uniqueness
with $r\in(1,d^{\prime})$ and $1{1}/{l}={1}/{r}+{1}/{d}-$ 1. Thus, we have
@@
Let $(\mathfrak{p},\mathfrak{q})\in{\mathscr{I}}$ be the pair as in Theorem 1.8. Suppose that $u^{(1)}$ and $u^{(2)}$ are two distributional solutions to the $2$D NSE
@@
and that $u^{(i)}\in L^{\mathfrak{q}^{\prime}}_{t}\dot{W}^{1,\mathfrak{p}^{\prime}}_{x}$, $u^{(i)}(t)\in L^{2,\infty}_{x}$ and ×$\nabla\times u^{(i)}:=\partial_{1}u_{2}^{(i)}-\partial_{2}u_{1}^{(i)}\in C_{t}{\mathcal{M}}_{x}$, $i=1,2$. Then $u^{(1)}=u^{(2)}$.
@@
Suppose that $F$ is a divergence-free vector field in $L^{2,\infty}({\mathbb{R}}^{2};{\mathbb{R}}^{2})$ and ×$\nabla\times F\in{\mathcal{M}}$. Then
@@
Let $G:=K_{\mathrm{BS}}*(\nabla\times F)\in L^{2,\infty}$ and $H:=F-G$. Then, $\mathord{{\rm div}}H=0$ and ×$\nabla\times H=0$ in the sense of distribution. This implies
@@
That is, $\rho$ satisfies the $2$D vorticity NSE.
Therefore, by virtue of Theorem 1.8, we obtain ×$\nabla\times u^{(1)}=\nabla\times u^{(2)}$, which along with Lemma 5.3 yields the desired uniqueness $u^{(1)}=u^{(2)}$. ∎
@@ 6. Nonlinear Markov processes
Then $b\in L^{\infty}([s,T];L^{d,\infty}_{x})$ and $\mathord{{\rm div}}\,b=0$. Since $(\mu^{s,\zeta})$ lie in the Krylov class, any curve of probability measures satisfying (6.2) also belongs to the Krylov class. Thus, in view of the well-posedness result in Proposition 3.2, the linear Fokker-Planck equation (6.1) admits a unique probabilistic solution satisfying (6.2), which verifies condition $(b)$.
@@ Appendix B Interpolation Inequalities
where $[u]_{\alpha}$ is the Hölder seminorm of $u$. Estimate (B.3) was first proved by Nirenberg in [43].
@@
$\fint_{A}f:=\frac{1}{|A|}\int f~{}\mbox{ and }~{}\|{{-}}f\|{{-}}_{L^{p}(A)}:=\left(\fint_{A}|f|^{p}\right)^{\frac{1}{p}}.$
@@
$\underline{u}=\left(\int_{{\mathbb{R}}^{d}}\varphi_{R}\right)^{-1}\int_{{\mathbb{R}}^{d}}u\varphi_{R}~{}~{}\mbox{ and }~{}~{}\bar{u}=u-\underline{u}.$
@@
    $\|{{-}}\bar{u}\|{{-}}_{L^{q}(B_{R})}\leqslant CR\|{{-}}\nabla u\|{{-}}_{L^{q}(B_{R})}.$ (B.7)
@@
    $\|{{-}}\bar{u}\|{{-}}_{L^{p}(B_{R})}\leqslant CR\|{{-}}\nabla u\|{{-}}_{L^{q}(B_{R})}.$ (B.8)
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)-Project-ID 317210226-SFB 1283. D. Zhang is also grateful for the NSFC grants (No. 12271352, 12322108) and Shanghai Frontiers Science Center of Modern Analysis. The research of G. Zhao is supported by the National Natural Science Foundation of China (Nos. 12288201, 12271352, 12201611).

### v2

@@ 2.3. Sobolev regularity estimate
[…]
@@
[…]
@@
[…]
@@ 3.1. Existence
${\leqslant}$ $C\sum_{i=1}^{2}\|b_{i}-b_{i}^{N}\|_{L^{q_{i}}_{t}L_{x}^{p_{i}}}+C\sum_{i=1}^{2}\|b^{N}_{i}-b^{n}_{i}\|_{L^{q_{i}}_{t}L_{x}^{p_{i}}}+\int_{0}^{T}{\mathbf{E}}_{{\mathbf{Q}}}\left|b^{N}(t,Y^{n}_{t})-b^{N}(t,Y_{t})\right|\text{\rm{d}}t.$
@@ 3.2. Uniqueness
where $\langle\cdot,\cdot\rangle$ denotes the integration over $\mathbb{R}^{d}$.
@@
$\langle\zeta,v(0)\rangle=\int_{0}^{T}\langle\rho(s),f(s)\rangle\text{\rm{d}}s,\quad\forall f\in C_{c}^{\infty}((0,T)\times{\mathbb{R}}^{d}),$
@@ 4.1. SDEs with supercritical drifts
Let $g$ be a non-negative smooth function on $0,\infty)$ such that $g^{\prime}(r)\geqslant 0$ for all $r\geqslant 0$, $g(r)=0$ if $0\leqslant r\leqslant 1/2$ and $g(r)=1$ if $r>1$. Let
@@
and for $1\leqslant i\leqslant d-1$,
$b_{i}(x)$= $-Nx_{i}r^{1-d}\partial_{x_{d}}H(x)$ (4.3)
@@
for all $x_{d}>0$ and $1\leqslant i\leqslant d-1$. Taking into account (4.4) and the anti-symmetry of $b$ about the $x_{d}$-axis, we obtain that $\mathord{{\rm div}}b(x)=0$ if $x\neq 0$.
@@
Since $0<1/(1+\alpha)<1/2$, we can choose $N\gg 1$ such that
@@
The above inclusion indicates that the solution paths starting from $x\in{\mathcal{C}}_{\kappa,2}$ stay in the larger cone ${\mathcal{C}}_{1}$ over a certain time interval $0,t_{x}]$, and return back to the small cone ${\mathcal{C}}_{\kappa}$ at time $t_{x}$.
@@
$\begin{aligned} V_{t}^{x}(\omega)-x_{d}=&\int_{0}^{t}b_{d}(\omega_{s})\text{\rm{d}}s=N(d-1)\int_{0}^{t}[V^{x}_{s}(\omega)+(B_{s}(\omega))_{d}]^{-\alpha}\text{\rm{d}}s\geqslant 0,\\ \widehat{\omega}_{t}-\hat{x}=&N\alpha\int_{0}^{t}[V^{x}_{s}(\omega)+(B_{s}(\omega))_{d}]^{-\alpha-1}\widehat{\omega}_{s}\text{\rm{d}}s+\widehat{B}_{t}(\omega).\end{aligned}$ (4.13)
@@
$\begin{aligned} \underline{V}_{t}^{x}(\omega):=&\left[x_{d}^{1+\alpha}+(1+\varepsilon)^{-\alpha}N(d-1)(\alpha+1)t\right]^{\frac{1}{1+\alpha}}\\ \leqslant&V_{t}^{x}(\omega)\leqslant\left[x_{d}^{1+\alpha}+(1-\varepsilon)^{-\alpha}N(d-1)(\alpha+1)t\right]^{\frac{1}{1+\alpha}}=:\overline{V}_{t}^{x}(\omega),\end{aligned}$ (4.15)
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.2. FPE with supercritical drifts
[…]
@@
where the last identity was due to the uniqueness assumption and the fact that the regular conditional probability ${\mathbb{P}}_{x}(\cdot|{\mathcal{F}}_{\sigma})(\omega^{\prime})$ is also a martingale solution to (SDE) starting from 0, for ${\mathbb{P}}_{x}$-a.s. $\omega^{\prime}$. Taking into account that the distribution of $(\sigma,\omega_{\cdot\wedge\sigma})$ is uniquely determined under ${\mathbb{P}}_{x}$, we thus infer that any one-dimensional marginal distribution of weak solutions to (SDE) satisfies the same identity (4.2), and thus, is unique for all $x\in{\mathbb{R}}^{d}$.
@@ 5.1. Existence
$\left\{\begin{aligned} \text{\rm{d}}X_{t}^{n}&=K^{n}(t,X_{t}^{n}-y)\rho^{n}(t)(\text{\rm{d}}y)+\sqrt{2}\text{\rm{d}}W_{t}\\ \rho^{n}(t)&=\mathrm{law}(X_{t}^{n}),~\rho^{n}(0)=\zeta\in{\mathcal{P}}({\mathbb{R}}^{d}).\end{aligned}\right$.
@@
$\int_{0}^{t}\langle\rho^{n}(s),K^{n}*\rho^{n}(s)\cdot\nabla\phi(s)\rangle\,ds\to\int_{0}^{t}\langle\rho(s),K*\rho(s)\cdot\nabla\phi(s)\rangle\,ds,\quad\phi\in C_{c}^{\infty}([0,T]\times{\mathbb{R}}^{d}),$
@@ 5.2. Uniqueness
with $r\in(1,d^{\prime})$ and 1/l=1/r+1/d1{1}/{l}={1}/{r}+{1}/{d}-1. Thus, we have
@@
Let $(\mathfrak{p},\mathfrak{q})\in{\mathscr{I}}$ be the pair as in Theorem 1.8. Suppose that $u^{(1)}$ and $u^{(2)}$ are two distributional solutions to the 22D NSE
@@
and that $u^{(i)}\in L^{\mathfrak{q}^{\prime}}_{t}\dot{W}^{1,\mathfrak{p}^{\prime}}_{x}$, $u^{(i)}(t)\in L^{2,\infty}_{x}$ and $\nabla\times u^{(i)}:=\partial_{1}u_{2}^{(i)}-\partial_{2}u_{1}^{(i)}\in C_{t}{\mathcal{M}}_{x}$, $i=1,2$. Then $u^{(1)}=u^{(2)}$.
@@
Suppose that $F$ is a divergence-free vector field in $L^{2,\infty}({\mathbb{R}}^{2};{\mathbb{R}}^{2})$ and $\nabla\times F\in{\mathcal{M}}$. Then
@@
Let $G:=K_{\mathrm{BS}}*(\nabla\times F)\in L^{2,\infty}$ and H:=FGH:=F-G. Then, $\mathord{{\rm div}}H=0$ and $\nabla\times H=0$ in the sense of distribution. This implies
@@
That is, $\rho$ satisfies the 22D vorticity NSE.
Therefore, by virtue of Theorem 1.8, we obtain $\nabla\times u^{(1)}=\nabla\times u^{(2)}$, which along with Lemma 5.3 yields the desired uniqueness $u^{(1)}=u^{(2)}$. ∎
@@ 6. Nonlinear Markov processes
[…]
@@ Appendix B Interpolation Inequalities
where $u]_{\alpha}$ is the Hölder seminorm of $u$. Estimate (B.3) was first proved by Nirenberg in [43].
@@
$\fint_{A}f:=\frac{1}{|A|}\int f~\mbox{ and }~\|\hskip-6.4pt{{-}}f\|\hskip-6.4pt{{-}}_{L^{p}(A)}:=\left(\fint_{A}|f|^{p}\right)^{\frac{1}{p}}.$
@@
[…]
@@
    $\|\hskip-6.4pt{{-}}\bar{u}\|\hskip-6.4pt{{-}}_{L^{q}(B_{R})}\leqslant CR\|\hskip-6.4pt{{-}}\nabla u\|\hskip-6.4pt{{-}}_{L^{q}(B_{R})}.$ (B.7)
@@
    $\|\hskip-6.4pt{{-}}\bar{u}\|\hskip-6.4pt{{-}}_{L^{p}(B_{R})}\leqslant CR\|\hskip-6.4pt{{-}}\nabla u\|\hskip-6.4pt{{-}}_{L^{q}(B_{R})}.$ (B.8)
@@ Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation)-Project-ID 317210226-SFB 1283. D. Zhang is also grateful for the NSFC grants (No. 12271352, 12322108) and Shanghai Frontiers Science Center of Modern Analysis. The research of G. Zhao is also supported by the NSFC grants (Nos. 12288201, 12271352, 12201611) and National Key Research and Development Program of China (No. 2024YFA1013503).